How do you find the x, y coordinates?
y+4x-8 , y=2x=10
2x-3y=-2 , y=-4x+24
2x-10y=-12 , 3x+y=4
-3x+6y=48 , -5x+6y=32
x=6y-4 , x=-2y+28

Answer by ishan sethi(3)

2x-3y=-2 or (2x-3y+2=0)*2
y=-4x+24 or 4x+y-24=0
4x-6y+4=0_(1)
4x+y-24=0_(2)
subtract 1 from 2
u will get 7y-28=0
therefore,y=4&x=5
LIKEWISE SOLVE OTHER PROBLEMS TOO.